03. d coolant circulating at a temperature of 93°C through the circulating (D A certain car engines cooling system. Assume that, in this normal condition, the coolant complete aluminum radiator and the 10.5-L internal cavities within the When a car overhea ex ) A certain car has 14-0 L of liquid fills the 3.5-L volume of the and coolant expand and a small reservoir connected reservoir if the system goes from 93°C to 105°C. ts, the radiator, engine, to the radiator catches any resultant coolant overflow Estimate how much coolant overflows to the reservoir ifthe as hollow shells of aluminum. The coefficient of volume Model the radiator and engine expansion for coolant isB 410 x 106/Co 04. more than 200 years ago, was a platinum bar with two The first real length standard, adopted very fin e marks separated by what was defined to be exactly one meter. If this standard was to be accurate to within 1.0 um how carefully would the trustees have needed to control the temperature? The coefficient of linear expansion is 9 x 1 0-6 / C。 SECTION II Q5 A 215-g sample of a substance is heated to 330 C and then plunged into a 105-g aluminum calorimeter cup containing 185 g of water and a 17-g glass thermometer at 10.5°C. The final temperature is 35.0°C. What is the specific heat of the substance? (Assume no water boils away.) Identify the substance.
Answer #1

MANGE IN VOLUME CaN BE CALcCULATED AS- ATOR dy,. (3.51) (75x10-5) (105-93%) : = 3.15 x10-31 IN CAVITIES 9.45×10-3 L Hus, THE NEH) voluMES FOR THE RADaTo« aNd dv, 3.5L+ 3.15x10-3 L 3.50315 L d½ 10.5 L +9.45 x 10-3 L с 10.50945 L = INTERNAL CAVITI ES ARE - . 3. 50 3 1 5 L + 10. 50945 L = 14.0126 L 八.. у裋74.: 4. ecA.JP, DETERMNE THE VoLuME ExPANSION OF THE .n CODLANT 0. 06888 L THE NEW VOLuME IS- 14. 06888 L CALCULATE THE AMOUNT OF cOOLANT THAT OVERfLOws To THE RESERVDIR C TOT TO T e 14.06888L14. 0126 L « o.o5628 L 56.28 ml4> THE EQuATION TO CONTROL THE ד EMPERATURE SOLVE FOR THUs THE TrUSTEES CON TROLLED THE TEMPERA- TURE TO AN ACCURACY OFİ+ o.1ร่า HEAT LOST 6Y SuBS TANCE HEAT WATER aND ALUMiNIum ANd Glass THERMDMETER SuestaNcE 330- 35 °c x (35-10.5%) (o.215 CSUB (330-35%) [(o.kes ) (4,86 z/ל.c) + (o.10%)(qooz/ל.C) 63. 425 Csus * 14.28] (24.5) 774.41 + 94.5 CSug ะ L774.41 +94.5+14.28 ] (24.5) 6 3.425